OpenStudy (anonymous):

evaluate the integral from x=1 to x=2 of (v^3 +4v^5)/v^2

OpenStudy (anonymous):

Is this your question? \[\int\limits_{1}^{2}\frac{v ^{3} +4v ^{5}}{v ^{2} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If so note common denominator: v^2. So you can rewrite: \[\frac{ v^{3} }{ {v^{2}} } + \frac{ v^{5} }{ v^{2} }\]

OpenStudy (anonymous):

This should give you v+v^3 So now you are integrating: \[\int\limits_{1}^{2} v+4v^{3}\]

OpenStudy (anonymous):

According to the properties of integrals: \[\int\limits_{1}^{2}v dv + \int\limits_{1}^{2}v ^{3}dv\]

OpenStudy (anonymous):

Do you understand so far?

OpenStudy (anonymous):

Sorry I left out the 4 that should be multiplying v^3.

OpenStudy (anonymous):

yeah it's making sense.

OpenStudy (anonymous):

Now we integrate both and add the result: Do you know how to integrate?

OpenStudy (anonymous):

kind of, i'm not very good at it

OpenStudy (anonymous):

Ok. \[\int\limits_{1}^{2} vdv = \frac{ 1 }{ 2 }(v^{1+1})\] with the range from 1 to 2 So this evaluates to, 1/2(2^2) - 1/2(1^2)

OpenStudy (anonymous):

First part would evaluate to 2-1/2 = 3/2 What i did was when you integrate a variable raised to a power you add 1 to the exponent and multiply the result by the reciprocal of the exponent: 1/2 Can you try the second part.

OpenStudy (anonymous):

Okay, ya i'll try it

OpenStudy (anonymous):

i got 30

OpenStudy (anonymous):

So what did you get when you integrate 4v^3? Did you get (4)*(1/4)*(v^4) ? which simplifies to v^4

OpenStudy (anonymous):

ohh, no i did that wrong then

OpenStudy (anonymous):

i figured it out. that would result in 15 and the total is 33/2

OpenStudy (anonymous):

Good. Remember the first value we got was 3/2 So 3/2+15 = ?

OpenStudy (anonymous):

Yup that is correct. Good job

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You are welcome!